### v1

# A Unified Data Augmentation Framework for Low-Resource Multi-Domain Dialogue Generation

Abstract
Current state-of-the-art dialogue systems heavily rely on extensive training datasets. However, challenges arise in domains where domain-specific training datasets are insufficient or entirely absent. To tackle this challenge, we propose a novel data Augmentation framework for Multi-Domain Dialogue Generation, referred to as AMD2G. The AMD2G framework consists of a data augmentation process and a two-stage training approach: domain-agnostic training and domain adaptation training. We posit that domain corpora are a blend of domain-agnostic and domain-specific features, with certain representation patterns shared among diverse domains. Domain-agnostic training aims to enable models to learn these common expressive patterns. To construct domain-agnostic dialogue corpora, we employ a de-domaining data processing technique used to remove domain-specific features. By mitigating the effects of domain-specific features, the model trained on the de-domained corpora can effectively learn common expression patterns in different domains. Subsequently, we adapt the learned domain-agnostic features to the target domain through domain adaptation training. We conduct experiments on Chinese dialogue datasets from five different domains and show that AMD2G achieves superior performance compared to both direct training on the target domain corpus and collective training on all five domain corpora. Our work underscores AMD2G as a viable alternative solution for low-resource multi-domain dialogue generation. Code and data associated with our work are available on GitHub repository.
Keywords:
dialogue generation domain adaptation low-resource.1 Introduction
The efficacy of established sequence-to-sequence methodologies in constructing dialogue systems has demonstrated remarkable success in previous research [39, 19, 20, 13]. More recently, the notable achievements of Large Language Models (LLMs), including Blender [37], Meena [1], ChatGPT [27], and GPT-4 [26], have prompted the research community to increasingly embrace generative models as the go-to approach. However, training these models often requires huge corpora. Unfortunately, the availability of domain-specific corpora remains notably limited across many domains, including medicine, finance, and military, due to concerns on security, copyright, and other constraints [6]. On the other hand, while the current LLMs exhibit excellent comprehension and generation capabilities, they often suffer from hallucinations in tasks with obvious domain characteristics and strong factuality [3, 21, 8], particularly in low- and zero-resource contexts. Hence, devising strategies to leverage corpora from disparate domains to enhance the performance and accuracy of the target domain remains a relevant pursuit in the era of LLMs.
Existing methods for cross-domain dialogue generation can be categorized into three groups [32]: i) Separate Pattern [43, 31, 47, 25, 14]; ii) Mixed Pattern [24, 45, 17, 9, 51, 23]; and iii) Shared-Private Pattern [54, 4, 46, 44, 2]. The Separate Pattern involves training the model separately for each domain, necessitating an adequate corpus for each domain. In contrast, the Mixed Pattern combines multi-domain datasets by prioritizing domain-agnostic features while disregarding domain-specific ones. However, models based on the Mixed Pattern may struggle to capture features from low-resource domains compared to high-resource ones. The Shared-Private Pattern extracts domain-agnostic and domain-specific features using shared and private modules, respectively, also relying on sufficient domain corpora. While these methods have demonstrated promising results, they all presuppose the availability of ample corpus data.
In the context of dialogue generation within low-resource domains, the effective leverage of resources from other domains holds significant importance. As illustrated in Figure 1, the domain corpora encompass both domain-agnostic and domain-specific information. For instance, in the film domain, the expression “Do you know the movie Avatar" domain comprises domain-specific information “movie Avatar” alongside the domain-agnostic fragment “Do you know the …”. Similarly, phrases like “song Love" and “Khoomei" carry domain-specific features of the music field. Notably, disparate domains exhibit shared expression patterns within their corpora. Leveraging these shared features provides an opportunity to enhance the performance of dialogue generation in low-resource domains.
Accordingly, we propose a simple yet effective data Augmentation method for Multi-Domain Dialogue Generation, termed AMD2G. As shown in Figure 2, we initially build a domain dictionary for each domain automatically. Subsequently, the domain corpora undergo de-domaining through the usage of domain dictionaries. Specifically, placeholders replace domain-specific keywords in the corpus identified in the domain dictionary.
The models are then fine-tuned on the combined de-domained corpora to learn common representation patterns across different domains. Finally, low-resource fine-tuning is conducted on the target domain dataset to acquire domain-specific features. We conduct experiments with AMD2G using Chinese conversation datasets from five distinct domains. This choice is motivated by the morphological simplicity of the Chinese language, characterized by few inflectional variations, aligning well with our de-domaining processing at the lexical level. We compare our proposed method with two baselines: direct training on the target domain corpus and training on all five domain corpora collectively. Our experimental findings demonstrate that the integration of AMD2G consistently enhances model performance across all five domains.
In summary, our contributions are as follows:
We introduce de-domaining for multi-domain dialogue generation datasets, a data augmentation technique that effectively reduces the impact of domain-specific features by extracting shared representations across domains.
We propose AMD2G, a simple yet effective alternative framework for the multi-domain dialogue generation task in low-resource settings.
We conduct experiments with AMD2G across five domains, demonstrating its superiority over direct training in the target domain and joint training across all five domains.
2 Methodology
As depicted in Figure 2, AMD2G primarily comprises two steps: data processing and training. In data processing, the domain corpus is de-domained by using the constructed domain dictionaries. In the training step, domain-agnostic training first allows models to learn shared patterns among multiple domains, while the domain adaptation phase enables models to capture domain-specific features.
2.1 Problem Formulation
An instance for one domain can be represented as , where = with utterances representing the context of the dialogue. Here, represents the -th utterance, and represents the corresponding response. Our goal is to build a corresponding dialogue generation system using corpora from other domains under the low-resource condition in the target domain. Please refer to the experimental settings for detailed settings of low resources (i.e., Section 3.4).
Note that AMD2G can be applied to models of both encoder-decoder and decoder-only structures. For models with an encoder-decoder structure, the encoder is responsible for encoding the dialogue history , and the decoder generates responses based on the encoded representation. For models with a decoder-only structure, we concatenate the dialogue history and responses into a consecutive sequence and perform sequence modeling by autoregression.
2.2 De-Domaining Data Processing
We observe that data from different domains exhibits shared representation patterns. Through de-domaining operations, corpora from different domains can be transformed into a unified space devoid of domain-specific features. This process mitigates the influence of domain-specific features, facilitating the learning of domain-independent features by models.
2.2.1 De-Domaining
Domain-specific corpora are de-domained based on the usage of domain dictionaries. Specifically, we replace all words or phrases present in the domain dictionary with designated placeholders, wherein all tokens within an involved phrase are substituted by a single placeholder. As a result of the de-domaining processing, the domain-specific data no longer retains its specific characteristics.
2.2.2 Dictionary Construction
We combine LLM-extracted terms with existing term banks to construct high-quality domain dictionaries tailored to specific domains. Initially, we employ TechGPT [36], a Chinese LLM, to extract domain entities as keywords. TechGPT has been enhanced for various information extraction tasks through the integration of domain knowledge graphs [35] facilitated by BELLE [52], which is specialized in extracting domain keywords. We apply the following prompt to TechGPT to extract entities from the domain corpora. $Context$ represents the context composed of dialogue history and response. $Domain$ represents the domain name.
In addition to keywords extracted by the LLM, we utilize existing terminology banks offered by Chinese input method providers, including QQPinyin111http://cdict.qq.pinyin.cn/v1, Baidu111https://shurufa.baidu.com/dict, and SougouPinyin111https://pinyin.sogou.com/dict/. We retrieve the terms specific to each domain from these sources and merge them into the respective domain’s dictionary. Table 1 presents the statistic overview of the dictionary, including its size, coverage ratio concerning the training set, and the number of replaced tokens. Notably, the domain dictionary’s coverage rate for the training corpus exceeds 95%, with a significant reduction in domain-specific terms.
Domain #Keyword #Cov #RToken Film 2463 100% 38680 Music 1427 100% 32305 Travel 1006 100% 33820 Medical 18749 99.80% 37065 E-comm 1384 95.10% 19602
Domain Uni Bi Tri Quad O2Music 77.67 32.41 15.73 7.80 O2Travel 67.01 25.42 12.30 5.62 O2Film 63.91 23.30 11.13 5.41 O2Ecomm 80.95 28.80 9.43 2.26 O2Medical 63.53 16.15 5.55 1.86
2.3 Domain-Agnostic Training and Domain Adaptation
We conduct the first stage of fine-tuning on a mixed domain-agnostic corpus excluding the target domain to learn domain-independent features. Models pay more attention to domain-independent features in the domain-agnostic training phase. The purpose of domain adaptation is to transfer domain-independent knowledge to the target domain, allowing models to learn domain-specific features. Subsequently, models are initialized with the weights from the domain-agnostic training stage and then fine-tuned on the low-resource target domain corpus.
2.4 Domain Similarity
We believe that the similarity between domains is an important factor for AMD2G. To explore the impact of different domains, we propose a simple domain similarity evaluation method. Specifically, we employ n-gram recall between domains as the similarity metric. Given that the n-gram sets of domains A and B are and respectively, the similarity score of domain A relative to domain B is:
The similarity of domain B to A can be computed in the same way. represents the degree of similarity between other domains and domain B. Theoretically, the greater the similarity between the mixed data and target domain, the greater the benefits of data augmentation to the target domain. Table 2 shows the average similarity scores between domains. Note that all results are based on the training set after removing domain words. Expression paradigms usually consist of more than two words, so similarity scores based on 2-gram and above can better show the degree of similarity between domains. An elevated similarity score signifies a greater overlap of paradigms between the target domain and other domains.
Model Corpus BLEU-1 BLEU-2 BLEU-3 BLEU-4 Rouge-L Dist-1 Dist-2 Embed A/E/G AVE PPL Film Transformer target 23.68 13.66 9.41 6.65 31.20 17.11 38.69 74.36 56.67 84.43 35.59 6.7480 mix 22.60 12.72 8.51 6.07 29.91 20.16 48.86 77.47 58.19 86.69 37.12 6.4009 AMD2G 29.67 16.96 11.92 8.70 33.52 22.92 51.18 78.64 59.94 86.87 40.03 6.2826 CPT target 26.18 14.65 10.05 7.24 31.32 25.34 56.69 78.07 58.81 86.67 39.50 4.4767 mix 26.15 14.90 9.86 7.01 32.14 26.10 58.39 78.28 59.38 86.92 39.91 3.9037 AMD2G 28.78 16.92 13.20 8.31 33.49 25.04 57.02 80.91 60.15 87.36 41.12 4.0225 GPT-2 target 8.92 4.53 2.70 1.55 21.11 6.46 18.47 75.87 57.68 87.83 28.51 7.7622 mix 10.74 5.20 3.05 1.73 24.67 10.11 25.49 75.11 56.74 86.77 29.96 7.6871 AMD2G 13.89 6.90 3.80 2.61 23.42 11.30 32.54 78.89 59.23 87.90 32.05 7.6083 BART target 26.52 14.81 10.17 7.50 31.25 29.39 60.04 76.93 58.41 86.75 40.18 3.9784 mix 25.64 15.23 11.06 8.46 32.73 29.47 61.73 78.32 59.93 86.39 40.90 3.4973 AMD2G 29.80 16.80 11.02 8.42 35.43 30.91 62.97 78.36 59.38 86.99 42.01 3.4282 Music Transformer target 32.49 18.20 12.53 9.45 34.04 13.65 30.98 77.17 61.17 86.58 37.63 4.8624 mix 36.37 21.23 13.98 9.72 39.19 17.12 39.06 81.93 65.84 88.75 41.32 4.2834 AMD2G 39.66 22.97 14.41 9.51 40.84 19.89 40.77 82.28 66.63 88.22 42.52 4.1936 CPT target 33.52 19.00 11.76 7.67 36.78 18.01 43.09 81.77 64.14 89.76 40.55 3.6496 mix 36.13 20.71 13.42 9.03 38.57 20.97 48.82 81.95 65.10 88.98 42.37 3.3757 AMD2G 37.41 21.24 14.23 10.05 39.80 21.50 50.68 82.63 65.89 89.64 43.31 3.2297 GPT-2 target 27.13 14.37 9.40 6.67 32.74 13.42 31.68 79.34 62.54 88.21 36.55 4.0748 mix 32.83 17.96 11.81 8.27 34.10 14.71 34.45 79.28 62.99 87.21 38.36 3.5407 AMD2G 35.98 19.49 13.83 9.53 35.00 18.81 44.34 81.99 63.31 88.40 41.07 3.3353 BART target 34.97 21.09 14.46 10.69 40.84 22.74 47.80 82.91 66.93 89.59 43.20 3.1210 mix 36.02 22.35 15.88 12.04 41.94 22.55 49.43 82.21 67.23 88.68 43.83 3.2139 AMD2G 38.25 23.99 15.81 14.62 43.46 23.53 53.89 82.59 68.09 89.69 45.39 3.0367 Travel Transformer target 34.63 27.17 22.57 19.43 47.26 11.55 25.24 80.44 70.83 86.92 42.60 2.3491 mix 36.03 28.52 23.65 20.58 46.85 12.34 31.90 83.90 72.15 90.22 44.61 2.3350 AMD2G 36.30 28.40 24.63 22.68 46.96 14.16 33.41 81.58 69.39 89.82 44.73 2.2389 CPT target 26.80 20.07 15.67 12.84 47.25 19.19 43.12 83.59 72.76 89.25 43.06 1.8730 mix 35.17 28.41 24.14 21.19 51.19 17.41 40.83 85.24 74.29 90.56 46.84 1.8273 AMD2G 36.67 28.73 24.41 21.91 51.93 17.07 39.85 86.48 74.96 91.29 47.33 1.8032 GPT-2 target 32.41 24.07 19.70 16.59 39.80 11.50 28.15 79.60 68.55 87.45 40.78 3.6998 mix 36.90 28.48 24.03 21.52 39.27 14.57 34.07 80.81 68.73 88.66 43.70 3.1826 AMD2G 37.65 30.09 25.88 23.18 44.60 12.07 27.44 83.11 71.77 89.41 44.52 3.0171 BART target 31.49 23.14 17.45 13.64 47.29 15.10 35.70 84.23 72.90 90.81 43.17 1.6796 mix 31.25 24.27 19.45 16.27 49.91 20.88 44.68 85.06 74.07 90.80 45.66 1.6497 AMD2G 36.23 28.93 24.18 20.72 51.57 19.85 41.85 84.75 74.34 91.25 47.37 1.6307 E-Commerce Transformer target 13.51 7.81 5.15 3.55 20.48 5.42 16.54 62.11 49.47 68.15 25.22 2.3847 mix 13.87 8.00 5.09 3.51 20.62 9.11 30.43 62.79 49.45 69.53 27.24 3.2139 AMD2G 14.92 9.71 6.49 5.97 21.84 9.37 31.70 64.67 51.54 69.17 28.54 2.3438 CPT target 13.25 8.28 5.70 3.97 23.07 12.02 36.63 63.42 51.04 70.80 28.82 2.1533 mix 15.55 9.67 6.91 5.33 22.29 12.88 40.92 63.83 50.41 71.87 29.97 2.5117 AMD2G 16.54 8.72 5.66 4.05 24.17 12.69 41.93 64.38 50.97 71.71 30.08 2.1168 GPT-2 target 7.21 3.69 2.17 1.37 13.70 3.01 9.91 60.58 46.49 71.09 21.92 2.2633 mix 5.52 2.64 1.62 1.16 10.50 3.60 11.01 58.52 44.01 69.00 20.76 2.2301 AMD2G 7.44 3.13 1.46 0.76 12.50 5.36 18.34 60.49 44.77 71.60 22.58 2.2175 BART target 14.94 9.04 6.10 4.18 21.95 13.23 38.38 62.88 50.43 70.25 29.14 2.0949 mix 15.60 10.56 8.23 6.99 22.30 13.78 40.75 62.39 49.92 71.05 30.16 1.9479 AMD2G 14.04 9.25 6.50 4.70 23.66 18.61 51.59 62.97 50.61 70.91 31.28 1.8750 Medical Transformer target 13.50 8.94 6.30 3.13 42.44 1.95 4.41 80.71 70.83 82.43 31.46 1.9193 mix 15.72 9.93 6.58 3.27 40.04 6.93 18.02 79.17 69.12 81.27 33.01 2.4432 AMD2G 18.50 11.33 7.24 3.98 37.81 9.39 25.10 78.39 67.17 81.52 34.04 1.8951 CPT target 14.03 9.29 6.46 2.49 40.85 13.68 29.23 78.31 68.46 81.43 34.42 1.9884 mix 19.88 12.50 8.32 4.62 39.37 13.75 36.04 78.69 67.54 82.21 36.29 1.6892 AMD2G 21.90 13.69 8.48 4.37 39.99 14.34 37.87 79.40 68.18 82.78 37.10 1.6861 GPT-2 target 19.13 11.12 6.82 3.08 35.33 4.09 10.46 80.67 70.22 83.00 32.39 6.0602 mix 23.70 14.16 8.90 4.33 36.42 4.53 12.10 81.14 70.92 83.32 33.95 5.7444 AMD2G 23.94 15.02 8.30 4.17 36.67 4.78 14.60 82.45 71.45 83.23 34.46 5.7559 BART target 13.05 8.84 6.35 3.01 42.69 8.22 15.05 79.97 70.55 82.21 32.99 1.9705 mix 16.67 11.12 7.79 4.43 42.77 11.72 27.45 79.74 69.65 82.79 35.41 1.9147 AMD2G 22.97 12.29 7.94 4.56 41.22 12.29 29.75 80.97 68.91 83.04 36.39 1.7770
3 Experiments
3.1 Datasets
In this paper, we experiment on Chinese dialogue generation datasets from five domains (i.e., Film, Music, Travel, Medical and E-commerce). Film, Music, and Travel are all from KdConv [55] datasets. KdConv is a Chinese multi-domain conversation dataset comprising 4.5K conversations from three domains: Film, Music, and Travel. It contains 86K utterances with an average turn number of 19.0. These conversations feature in-depth discussions and natural transitions between multiple topics. The Film, Music, and Travel domains contain 1,500 training sets, 150 validation sets, and 150 test sets respectively. MedDG is a large-scale Chinese medical dialogue dataset, which contains 14,864 training sets, 2,000 validation sets, and 1,000 test sets respectively [18]. E-commerce is a large-scale e-commerce conversation dialogue dataset, containing 500,000 positive training examples, 1,000 validation sets and 1,000 test sets respectively [53]. Similarly, we randomly selected 2,000 examples as the training set. In order to meet the low resource setting, we extract 2000 conversations as the training set for MedDG and E-commerce. It is worth noting that we do not use the knowledge base.
3.2 Models
To evaluate the effectiveness and robustness of our proposed method, we apply AMD2G to different types of models, including encoder-decoder and decoder-only structures. For the encoder-decoder structure models, we employ the basic Transformer [42] structure as well as Chinese pre-trained language models such as CPT [40] and the Chinese version of BART111https://huggingface.co/fnlp/bart-base-chinese [10]. The decoder-only structure model used in our experiment is the Chinese version of the pre-trained model GPT-2111https://huggingface.co/uer/gpt2-chinese-cluecorpussmall [33].
3.3 Baselines
For each model, we first compare AMD2G with training only on the original domain training set and training on the mixed training set of all domains. Besides, we compared AMD2G with two other multi-domain transfer methods: TS-NET [29] and DA-NET [23]. TS-NET uses a teacher-student network mechanism to transfer knowledge from other domains to the target domain, while DA-NET uses domain attention to realize knowledge transfer. We apply different methods to two types of models. One is the pre-trained Chinese model, BART, and the other is a language model that adopts the architecture of gated recurrent neural networks (GRU), a variant of RNN [5].
3.4 Implementation Details
We implement our model and baselines using the Huggingface library and train baselines on a server with RTX A6000 GPU (48G). We consider at most 10 turns of dialogue context and 50 words for each utterance. The batch size is 32, the minimum decoding length is set to 10, the maximum decoding length is set to 128, the warmup steps are 1,000, and the initial learning rate is 5e-5. We use the AdamW [22] optimizer to update model parameters. The beam size and length penalty coefficient are set to 6 and 1.0 respectively. The random seed of the sampled data is set to 12345. The random seed of the training process is set to 12345. We use GLoVe [30] to train 300-dimensional word vectors based on the training set for evaluation. The values of hyperparameters described above are all fixed using the validation set. We explore more low-resource scenarios including the ratio of [5%,10%,20%,30%,40%] of the target corpus.
3.5 Evaluation metrics
We follow the previous studies [20, 12, 16, 49, 19], using both automatic and human evaluations to assess the performance of models. Automatic evaluations include BLEU [50], ROUGE-L [15], Dist [11], and Embedded metrics. Perplexity is also measured as an additional metric. For human evaluation, we request annotators to score the generated responses with respect to three aspects: fluency, diversity, and relevance. Each dimension is divided into three levels: 0, 1, and 2. In terms of fluency, 0 means not fluency, 1 means average fluency, and 2 means very fluency. Other evaluation dimensions are similar to fluency. After collecting the assessments from annotators, the final score is the average of all samples. Note that we use an improved version of BLEU that is more in line with human evaluation, and the calculated score will be lower than the original BLEU [28].
4 Results and Analysis
Model Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 Rouge-L Dist-1 Dist-2 Embed A/E/G AVE PPL Film BART TS-NET 27.75 14.43 9.76 6.55 33.94 28.45 61.44 78.22 58.42 85.77 40.47 5.6721 AMD2G 29.80 16.80 11.02 8.42 35.43 30.91 62.97 78.36 59.38 86.99 42.01 3.4282 GRU DA-NET 15.55 11.01 7.38 6.33 22.44 11.07 23.08 67.43 54.76 83.55 30.26 25.6935 AMD2G 16.70 13.52 8.43 8.01 25.38 12.61 25.78 69.35 56.05 84.42 32.02 21.3866 Travel BART TS-NET 37.22 20.99 15.22 12.68 40.87 22.45 51.09 80.67 66.53 87.02 43.47 5.0060 AMD2G 38.25 23.99 15.81 14.62 43.46 23.53 53.89 82.59 68.09 89.69 45.39 3.0367 GRU DA-NET 17.86 14.55 10.22 8.55 27.69 14.58 27.82 71.84 56.62 85.33 33.51 12.6600 AMD2G 19.22 15.62 12.64 10.33 27.71 16.20 28.44 73.05 57.22 85.44 34.59 10.3700 E-Commerce BART TS-NET 13.69 10.04 6.72 4.55 21.79 16.93 29.02 60.44 49.79 69.94 28.29 1.9044 AMD2G 14.04 9.25 6.50 4.70 23.66 18.61 51.59 62.97 50.61 70.91 31.28 1.8750 GRU DA-NET 9.33 5.66 3.78 2.28 16.64 4.57 15.77 60.54 48.72 66.12 23.34 7.7743 AMD2G 11.62 6.44 4.44 3.22 20.09 5.67 17.66 60.62 49.90 67.03 24.67 7.1090
4.1 Overall Results
Table 3 reports the experimental results of AMD2G in five domains. Compared with training directly on the target domain training set, AMD2G demonstrates absolute advantages in five domains. Specifically, the average performance of the four models has improved by 1.85% in the e-commerce domain, 2.69% in the medical domain, 2.86% in the film domain, 3.59% in the music domain, and 3.58% in the travel domain, compared with training directly in the target domain. Figure 3 shows that the performance of some models using 30% target domain training corpus has achieved competitive performance based on AMD2G framework compared to training on the target domain training set. When using 40% target domain training corpus, the performance based on AMD2G is equivalent to training directly on the target domain training set. These results fully demonstrate the effectiveness of AMD2G.
According to Table 2, even though the two domains are quite different, they still share some expression paradigms, which accounts for why the AMD2G framework is effective. The first stage of the AMD2G framework allows the model to learn shared paradigms between different domains and the second stage allows the model to learn domain-specific features, which can maximize the learning of domain-agnostic features and reduce the mutual influence between domain features. Compared with the performance of training when mixing all domain corpora, the performance of models based on AMD2G exhibits absolute advantages, which confirms that features between domains can interfere with each other and has a negative impact on model performance. We will discuss in detail the impact of domain similarity on model performance in the next section. Table 4 reports the performance comparison of AMD2G and other baselines. We can observe that AMD2G has certain performance advantages compared to TS-NET and DA-NET. TS-NET uses a distillation mechanism to transfer domain knowledge, which relies on a large amount of target domain data. Low resource settings will severely impact distillation results. DA-NET utilizes a dynamic attention mechanism for multi-domain feature fusion. While it preserves domain-specific knowledge, it tends to overlook domain-independent features, thereby hindering its ability to effectively utilize features from other domains. The AMD2G adopts a two-stage strategy to retain domain-agnostic and domain-specific features, and domain-independent features can be adapted to the target domain and achieve data enhancement effects.
4.2 Impact of Domain Similarity
Domain similarity is a key impact factor for model performance. We perform further experiments to analyze the impact of similarity between domains on model performance. The last one in Figure 3 shows the distribution of domain similarity based on n-gram and the average performance gain of models based on AMD2G. We find that the similarity based on 1-gram does not accurately reflect the similarity relationship between domains because it is greatly affected by the unified placeholder. In fact, scores based on more than 2-gram can better reflect the similar relationship between domains, because the common expression paradigm is based on more than two words. According to the 4-gram similarity score, the data enhancement effect based on the AMD2G is more obvious for domains with high similarity scores to a certain extent. The more similar the domains are, the more helpful the knowledge provided by other domains will be to the target domain. An exception is the medical domain. Although the similarity score based on 4-gram is low, models based on the AMD2G framework have achieved a certain degree of gain. The domain characteristics of the medical domain are relatively obvious. After domain-free operations, the alignment of expression paradigms in different domains can be achieved to a certain extent. This can be derived from the similarity scores based on n-grams. This is why the AMD2G framework can work in the medical domain.
4.3 Impact of Dataset Size
In order to explore the impact of data set size, we conduct experiments on 5%, 10%, 20%, 30%, 40%, and 100% of the training set, respectively. Note that 100% of the training set refers to the total amount of data under low resources. Figure 3 reports the changing trends of model performance with data set size in five domains. We can observe that the average performance (i.e., AVE) basically increases gradually as the data size increases, and the PPL decreases as the data set size increases. Models are more sensitive to the size of the data set under low resource conditions. Even a small increase in the data set will make the performance of models increase. The increase in the data set will allow models to learn more different examples, making models cover more test examples.
4.4 Human Evaluation Results
Model Corpus E-Commerce Film Travel Fluency Relevance Diversity Fluency Relevance Diversity Fluency Relevance Diversity BART target 0.377 0.833 0.062 0.579 0.875 0.083 1.034 0.920 0.116 mix 0.464 1.025 0.085 0.662 0.965 0.098 1.142 0.946 0.104 AMD2G 0.522 1.150 0.113 0.784 1.110 0.115 1.206 1.012 0.133 GPT-2 target 0.311 0.753 0.065 0.466 0.782 0.076 0.972 0.938 0.096 mix 0.472 0.975 0.086 0.673 0.950 0.120 1.133 0.955 0.110 AMD2G 0.514 1.040 0.082 0.762 1.132 0.112 1.174 1.102 0.128
Model Method E-Commerce Film Travel Fluency Relevance Diversity Fluency Relevance Diversity Fluency Relevance Diversity BART TS-NET 0.456 0.814 0.065 0.677 1.067 0.108 1.110 0.864 0.105 AMD2G 0.522 1.150 0.113 0.784 1.110 0.115 1.206 1.012 0.133 GRU DA-NET 0.237 0.755 0.052 0.508 0.833 0.117 0.456 0.867 0.088 AMD2G 0.307 0.774 0.069 0.553 0.912 0.099 0.542 0.955 0.086
We conduct the human evaluation of BART and GPT-2 on three domains to further confirm the effectiveness of AMD2G. To evaluate the consistency of the results assessed by annotators, we employ Pearson’s correlation coefficient [38]. This coefficient is 0.25 on diversity, 0.68 on relevance, and 0.77 on fluency, with p < 0.0001 and below 0.001, which demonstrates high correlation and agreement. The results of the human evaluation are shown in Table 5. Compared to training directly in the target domain and on mixed corpora, models based on the AMD2G enjoy a significant advantage in relevance and diversity. Specifically, models based on AMD2G enjoy an average advantage of 20.40% in fluency, 24.1% in relevance, and 3.1% in diversity in three domains compared with models training directly in the target domain. Compared with models training on mixed corpora, models based on AMD2G enjoy an average advantage of 6.9% in fluency, 10.2% in relevance, and 1.3% in diversity in three domains. The experimental results show the effectiveness of the AMD2G framework. AMD2G can effectively reduce the mutual interference of domain features and strengthen the learning of domain-agnostic features. Table 6 reports the performance comparison of AMD2G and other baselines. The human results show that the AMD2G framework still outperforms the TS-NET and DA-NET. Specifically, the model’s average performance, based on AMD2G, is 9.8% higher than TS-NET and 4.3% higher than DA-NET.
5 Related Work
Dialogue systems are used as intelligent agents in various domains due to their ability to generate fluent and natural responses. Cross-domain learning refers to the technology of transferring knowledge from other domains to the target domain. The initial approach is to blend all domain corpus to learn domain-independent features. One of the drawbacks of this approach is the lack of domain-specific knowledge. Wu et al. [45] propose to use a global-to-local pointer mechanism search technique for external knowledge to enhance the domain-specific knowledge awareness of models. He et al. [7] employs pairwise similarity to distill contextually unrelated KB records to improve the quality of domain knowledge. Xie et al. [48] integrates domain-specific knowledge in the form of text-to-text format based on T5 [34]. Ma et al. [23] propose a domain attention module with distributional signatures of dialogue corpus to capture domain-specific knowledge. These methods may also lead to a long-tail distribution of domain data, making models trained severely biased under low-resource conditions. Another line of research is to train separate models for each domain focusing on domain-specific features. Madotto et al. [24] learns domain-specific features by combining external knowledge through a memory network [41]. [31] proposes to employ two-step retrieval and attention mechanisms to improve the quality of domain-specific features. Wu et al. [47] proposes to continue pre-training on domain corpus to adapt the language model to a specific domain. The shared-private framework, which combines the advantages of the above two, is a better choice. Zhong et al. [54] uses global modules to share parameters and local modules to learn domain-specific features. Wu et al. [46] allows domain-specific features to interact through shared-private mechanism. Bang et al. [2] learns task-related features by adding adapters for each task.
6 Conclusion
We propose a simple and effective data augmentation framework AMD2G for multi-domain low-resource dialogue generation. The domain characteristics of the corpus can be removed through domain dictionaries constructed by LLMs. Models trained on domain-independent corpus can reduce the interference of different domain features when models learn domain-independent features. Domain adaptation training can adapt the learned domain-independent features to the target domain. Experiments on four models in five domains demonstrate the effectiveness of the AMD2G framework. Compared with other baselines, the AMD2G framework has obvious advantages. AMD2G provides an alternative solution for low-resource multi-domain dialogue generation.
Acknowledgement
We would like to thank reviewers for their constructive comments. The project is supported by the National Natural Science Foundation of China (62172086, 62272092) and DFG (grant SCHU 2246/14-1). The project is also supported by China Scholarship Council.